# 3DGAA: Realistic and Robust 3D Gaussian-based Adversarial Attack for Autonomous Driving

### v1

@@ 4.3 Effectiveness of the Physical Filtering Module
Table 2: Filtering strategy comparison across different metrics.
@@
4.4 Effect of Shape Loss Design
As shown in Table 3, adding the shape loss slightly improves LPIPS (from 0.5516 to 0.5503) and SSIM (from 0.0141 to 0.0146), indicating better perceptual quality and structural preservation. The PSNR value remains similar, with only a minor drop (0.5884 to 0.5881), reflecting a shift from pixel-level perturbation to more physically realistic adjustments.
These results show that incorporating $\mathcal{L}_{shape}$ helps balance attack robustness and physical realism, aligning with our dual-objective design.
Table 3: Ablation study of the shape loss component.
@@ 4.5 Selective Dimension Optimization
As shown in Figure 6(a), optimizing geometric parameters (position $x$, scale $s$) leads to rapid increases in Label Confusion Rate (LCR), achieving strong attack effectiveness within 30 epochs. In contrast, optimizing texture parameters (opacity $\alpha$, color $c$) requires more iterations but ultimately reaches comparable or higher LCR. This indicates that geometry is preferable for fast convergence, while texture offers greater long-term attack potential. A detailed dimension-wise analysis in Figure 6(b) further reveals that position $x$ and scale $s$ are the most sensitive, followed by opacity $\alpha$ and color $c$, while rotation $q$ contributes minimally due to its global and non-linear influence. These findings support our selective optimization strategy introduced in Section 3.6, allowing practitioners to prioritize dimensions based on computational budget or deployment constraints. Detailed explanation is given in Appendix B.3
(a) Grouped optimization
(b) Single-dimension optimization
Figure 7: Physical-world adversarial attack under varying lighting: (a) ID, (b) IN, (c) OD, (d) ON. Row 1: vanilla, Row 2: w/o Aug., Row 3: w/ Aug. Only the fully augmented adversarial object consistently evades detection.
@@ 4.6 Transferability Across Models and Tasks
We evaluate the transferability of 3DGAA-generated adversarial objects across different perception models, including object detectors with diverse architectures and semantic segmentation models.
Cross-Detector Generalization. Table 4 reports the LCR performance of adversarial objects optimized on one detector and tested on others. Perturbations generated using two-stage detectors (Faster R-CNN and Mask R-CNN) exhibit stronger transferability, achieving LCR values as high as 6.180 on YOLOv3. In particular, attacks optimized on F-RCNN retain effectiveness on SSD (LCR = 2.193), suggesting that 3DGS-based adversarial modifications can partially generalize across detector families.
These results indicate that 3DGAA’s perturbations are not overfitted to specific model internals, and instead capture transferable vulnerabilities in perception pipelines, particularly benefiting from multi-dimensional optimization across geometry and texture.
Table 4: LCR ($\uparrow$) for adversarial objects optimized on one detector and tested across others.
@@
Semantic Segmentation Attack. We further test 3DGAA’s generalization to segmentation models by evaluating LCR on DeepLabV3 and FCN with ResNet backbones. As shown in Table 5, the attacks transfer reasonably well to segmentation settings, despite being optimized for detection. Interestingly, deeper backbones (ResNet-101) appear more vulnerable than ResNet-50 variants, suggesting that network depth may increase susceptibility to 3D physical perturbations.
Table 5: LCR ($\uparrow$) of segmentation models under 3DGAA attack.
@@ 4.7 Physical Realization
To verify the real-world effectiveness of 3DGAA, we conduct physical deployment experiments using miniature 3D-printed vehicles under four lighting conditions: indoor daylight (ID), indoor night (IN), outdoor daylight (OD), and outdoor night (ON). As illustrated in Figure 7, we compare three texture variants: (1) vanilla (non-adversarial), (2) 3DGAA without physical augmentation, and (3) 3DGAA with augmentation. The vanilla object is reliably detected across all conditions. In contrast, 3DGAA w/o Aug. only slightly reduces detection confidence, particularly under strong lighting. The full 3DGAA w/ Aug. achieves consistent detection failures or misclassifications across all four scenarios. These results demonstrate the critical role of the physical augmentation module in bridging the sim-to-real gap.
@@
Table 6: Physical-world AP@0.5 ($\downarrow$) comparison of attack methods.
@@ 5 Conclusion
We present 3DGAA, a novel adversarial object generation framework that jointly optimizes geometry and appearance in the 3D Gaussian Splatting space. Unlike prior methods, 3DGAA enables expressive, multi-view-consistent adversarial robustness. Furthermore, we propose a Physical Filtering Module for physical realism and a Physical Augmentation Module for simulating environmental to develop real-world deployment. Extensive experiments show that 3DGAA significantly degrades detection performance (mAP $\downarrow$ to 7.38%) while maintaining strong physical realism across detectors and environments. It also generalizes across viewpoints, model architectures, and segmentation tasks, validating 3DGAA as a robust and transferable pipeline for evaluating safety-critical perception systems.
We acknowledge that adversarial attacks can raise ethical and security concerns. A detailed discussion of potential societal risks and mitigation strategies is included in Appendix C.4.

### v2

@@ 4.3 Effectiveness of the Physical Filtering Module
Table 2: Filtering strategy comparison across different metrics. TP+SD achieves the best performance on physical realism and artifacts removal (AR).
@@
Table 3: Ablation study of the shape loss component. $\mathcal{L}_{adv}+\mathcal{L}_{shape}$ achieves the best physical realism.
@@
(a) Grouped optimization.
(b) Single-dimension optimization.
Figure 7: Physical-world adversarial attack under varying lighting. Column: ID, IN, OD, ON. Row 1: vanilla, Row 2: w/o Aug., Row 3: w/ Aug. Only the fully augmented adversarial object consistently evades detection.
4.4 Effect of Shape Loss Design
As shown in Table 3, adding the shape loss slightly improves LPIPS (0.5516 to 0.5503) and SSIM (0.0141 to 0.0146), indicating better perceptual quality and structural preservation. The PSNR value remains similar, with only a minor drop, reflecting a shift from pixel-level perturbation to more physically realistic adjustments. These results show that incorporating $\mathcal{L}_{shape}$ helps balance attack robustness and physical realism, aligning with our dual-objective design.
@@ 4.5 Selective Dimension Optimization
As shown in Figure 6(a), optimizing geometric parameters (position $x$, scale $s$) leads to rapid increases in Log Confidence Reduction (LCR), achieving strong attack effectiveness within 30 epochs. In contrast, optimizing texture parameters (opacity $\alpha$, color $c$) requires more iterations but ultimately reaches comparable or higher LCR. This suggests a time-performance trade-off: geometry enables fast convergence, while texture offers stronger long-term perturbation due to its direct effect on visual semantics.
A detailed dimension-wise analysis in Figure 6(b) further reveals that position $x$ and scale $s$ are the most sensitive parameters, contributing significantly to early-stage attack progress. Opacity $\alpha$ and color $c$ exhibit moderate sensitivity, as they influence visibility and texture realism, respectively. Rotation $q$, despite affecting global geometry, shows minimal impact, due to its non-local and redundant nature in Gaussian-based representations. These findings validate our design of selective dimension optimization (Section 3.6), which allow practitioners to balance computational efficiency and physical feasibility in real-world deployments. Additional explanations and visual examples are provided in Appendix B.3.
@@ 4.6 Transferability Across Models and Tasks
We evaluate the transferability of 3DGAA adversarial objects across different perception models, including detectors with diverse architectures and segmentation models.
Cross-Detector Generalization. Table 4 reports the LCR performance of adversarial objects optimized on one detector and tested on others. Perturbations generated from two-stage detectors (Faster R-CNN and Mask R-CNN) demonstrate strong transferability, reaching LCR = 6.180 on YOLOv3 and maintaining performance on SSD (e.g., F-RCNN $\rightarrow$ SSD yields LCR = 2.193). These results suggest that 3DGAA captures transferable vulnerabilities across architectures through multi-dimensional optimization, rather than overfitting to a specific detector.
Table 4: LCR ($\uparrow$) for adversarial objects optimized on one detector and tested across others. 3DGAA achieves strong cross-detector generalization.
@@
Segmentation Attack. We further test 3DGAA’s generalization to segmentation models by evaluating LCR on DeepLabV3 and FCN with ResNet backbones. As shown in Table 5, the attacks transfer reasonably well to segmentation settings, despite being optimized for detection. Interestingly, deeper backbones (ResNet-101) appear more vulnerable than ResNet-50 variants, suggesting that network depth may increase susceptibility to 3D physical perturbations.
Table 5: LCR ($\uparrow$) of segmentation models under 3DGAA attack. 3DGAA achieves strong performance on segmentation models.
@@ 4.7 Physical Realization
To verify the real-world effectiveness of 3DGAA, we conduct physical deployment experiments using miniature 3D-printed vehicles under four lighting conditions: indoor daylight (ID), indoor night (IN), outdoor daylight (OD), and outdoor night (ON). As illustrated in Figure 7, we compare three texture variants: (1) vanilla (non-adversarial), (2) 3DGAA without physical augmentation, and (3) 3DGAA with augmentation. The vanilla object is reliably detected across all conditions. In contrast, 3DGAA w/o Aug. only slightly reduces detection confidence, particularly under lighting. The full 3DGAA w/ Aug. achieves consistent detection failures or misclassifications across all four scenarios. These results demonstrate the critical role of the physical augmentation module in bridging the sim-to-real gap.
@@
Table 6: Physical-world AP@0.5 ($\downarrow$) comparison of attack methods. 3DGAA achieves superior physical adversarial effectiveness.
@@ 5 Conclusion
We present 3DGAA, a novel adversarial object generation framework that jointly optimizes geometry and appearance in the 3D Gaussian space. Unlike prior methods, 3DGAA enables expressive, multi-view-consistent adversarial robustness. Furthermore, we propose a Physical Filtering Module for physical realism and a Physical Augmentation Module for simulating environmental to develop real-world deployment. Extensive experiments show that 3DGAA significantly degrades detection performance (mAP $\downarrow$ to 7.38%) while maintaining strong physical realism across detectors and environments. It also generalizes across viewpoints, model architectures, and segmentation tasks, validating 3DGAA as a robust and transferable pipeline for evaluating safety-critical perception systems, especially in autonomous driving scenarios.
A detailed discussion of security concerns and potential societal risks and mitigation strategies is included in Appendix C.4.